OpenStudy (anonymous):

PLEASE HELP What transformation has changed the parent function f(x) = log3x to its new appearance shown in the graph below

OpenStudy (mathmate):

Hint: When transforming f(x)->g(x) in the form g(x)=f(x-h)+k, then g(x) is translated from f(x) to the right by h, and up by k. We know that when \(f(x)=log_3(x)\), there is vertical asymptote at x=0 towards negative infinity, and that f(1)=0, and f(3)=1. Observe the graph of g(x) to see how g(x) was transformed from f(x), and work out the transformation accordingly.

OpenStudy (anonymous):

@mathmate i got f(x) – 2

OpenStudy (anonymous):

is it right? @mathmate

OpenStudy (unklerhaukus):

f(x)-2, would shift the graph down the vertical axis, but you want it to shift to the left

OpenStudy (anonymous):

im really bad at this so would it be f(x-2)

OpenStudy (unklerhaukus):

that is a shift to the right

OpenStudy (anonymous):

oh ok so it would be positive? f(x) = log3^(x+2)

OpenStudy (unklerhaukus):

i think you mean g(x) = f(x+2) = log_3 (x+2)

OpenStudy (anonymous):

yes would that be correct

OpenStudy (unklerhaukus):

yes!

OpenStudy (anonymous):

Thank you :)

OpenStudy (unklerhaukus):

when x approaches -2, the log has a vertical asymptote